### v1

# Investigating generalization capabilities of neural networks by means of loss landscapes and Hessian analysis

Abstract
This paper studies generalization capabilities of neural networks (NNs) using new and improved PyTorch library Loss Landscape Analysis (LLA). LLA facilitates visualization and analysis of loss landscapes along with the properties of NN Hessian. Different approaches to NN loss landscape plotting are discussed with particular focus on normalization techniques showing that conventional methods cannot always ensure correct visualization when batch normalization layers are present in NN architecture. The use of Hessian axes is shown to be able to mitigate this effect, and methods for choosing Hessian axes are proposed. In addition, spectra of Hessian eigendecomposition are studied and it is shown that typical spectra exist for a wide range of NNs. This allows to propose quantitative criteria for Hessian analysis that can be applied to evaluate NN’s performance and assess its generalization capabilities. Generalization experiments are conducted using ImageNet-1K pre-trained models along with several models trained as part of this study. The experiment include training models on one dataset and testing on another one to maximize experiment similarity to model performance “in the Wild”. It is shown that when datasets change, the changes in criteria correlate with the changes in accuracy, making the proposed criteria a computationally efficient estimate of generalization ability, which is especially useful for extremely large datasets.
Keywords: Neural networks, loss landscape, Hessian analysis, vizualization, generalization.
1 Introduction
Neural networks (NNs) have become indispensable in many fields of science and technology, including computer vision, robotics, cybersecurity, medicine, and others. NNs evolved significantly from small proof-of-concept models to large-scale industrial applications where processing enormous amounts of new data is necessary. Modern NNs are often trained on extremely large datasets under the assumption that this would result in good generalization to cases “in the Wild” [1, 2]. However, there is still very little research into optimal conditions that can ensure good generalization of NNs.
This increases the demand for analysis methods that can provide additional information about NN performance other than primitive “accuracy” calculation. In this paper we study such method which is associated with plotting and analyzing loss landscapes of NNs to assess NN stability and generalization capabilities [3, 4]. It allows to explore loss variation for specific models under different conditions which has been shown to correlate with training quality and generalization capability. The results can be visualized in a variety of ways and a research question regarding the optimal method is still open.
In this paper we present a new PyTorch [5] library called Loss Landscape Analysis (LLA) [6] which combines various approached to loss landscape plotting with analysis methods that can provide qualitative and quantitative assessment metrics of NN performance. We significantly extend an older library [7] by adding new analysis modes, different evaluation axes support, new weight update methods, etc. LLA also incorporates modern techniques for Hessian analysis [8, 9].
Whereas the loss landscape research field is not new, there are still very few criteria that can provide quantitative analysis metrics. To bridge this gap, we investigate the behavior of loss landscapes and Hessians of different NNs and formulate criteria for their analysis. We then conduct extensive generalization experiments to assess the viability of the proposed criteria. We also show that conventional loss landscape analysis techniques may yield incorrect results when applied to modern networks, showing that additional research in this field is needed. We primarily focus on NNs trained to solve classification tasks using supervised learning [10], and all results reported in this paper are obtained using LLA.
The rest of the paper is organized as follows: Section 2 discusses loss landscape analysis methodology and highlights existing problems, Section 3 studies Hessians of conventional NNs in different regimes and proposes criteria for Hessian analysis, Section 4 summarizes generalization experiment results and investigates the applicability of the proposed analysis criteria, and Section 5 concludes the paper.
2 Loss landscape analysis methodology
Loss landscapes are obtained by varying model weights and calculating corresponding values of a specific loss function. Since weights’ dimension is extremely large, several (typically one or two) direction vectors are chosen to plot the landscapes, with vectors’ dimensions matching the dimensions of weights. These vectors commonly have all values randomly taken from uniform or normal distribution [3, 7], and they are referred to as random directions. Figure 1 shows loss landscapes for convolutional NN (CNN) ResNet18 [11] and Visual Transformer (VIT) [12] plotted along random directions.
However, one should be mindful of the scale at which loss landscapes are plotted, since this can impact the way its surface is perceived. The most common weight update equation used for loss calculation is
$L=f\left(w+a\cdot d_{\text{1}}+b\cdot d_{\text{2}}\right),$ (1)
where w are the original weights, and a and b are coefficients corresponding to direction vectors d1 and d2, respectively. Loss landscape is plotted by varying a and b in steps and recording the resulting loss value. It should be noted that the choice of the range for a and b, which defines how much the weights would change, is somewhat arbitrary. In LLA we follow the procedure proposed in [7] that uses 40 steps in [-20,20] range for both coefficients, with -20 corresponding to 0th step and 20 corresponding to 40th step. This also places the original unmodified weights on [20,20] point in the middle of the plot, as can be seen in Figure 1. This range choice has allowed us to observe unexpected behavior of NNs with batch normalization (BN) discussed in Section 2.3.
Figure 1: Loss landscapes of (left) ResNet18, and (right) VIT-small plotted along random axes.
2.1 Choosing direction vectors for landscape plotting
At first glance the idea to use random directions in high-dimensional space might not seem very promising, since one might expect to get very different results depending on the choice of directions. However, the landscapes plotted along random directions are not as “random” as one might think. It was previously shown that the number of non-degenerate directions in parameter space decreases drastically during training [13]. Therefore, there is only a handful of directions that allows to obtain significantly different results. That being said, finding multiple directions to obtain different perspectives on the landscape might indeed be challenging, which is one of the shortcomings of random axes method [14].
Fortunately, there are several deterministic methods to choose the directions. The first one has to do with the underlying geometry of the parameter space, namely the Hessian of the neural network weight matrix [14], which will be discussed in detail in Section 3.1. The information about the landscape one obtains with Hessian axes is much richer, as will be shown in the next subsection. Another way is to use optimizer parameter axes as vectors. For models trained with Adam optimizer [15], these parameter vectors are moment vectors that influence the way model weights are altered during training. This allows to obtain significantly different loss landscapes analyzing which is out of the scope of this paper. The options to use Hessian and Adam axes for loss landscape plotting is implemented in LLA library.
2.2 Normalization of direction vectors
Normalization of direction vectors is often necessary since random values between 0 and 1 are added to weights of undefined scale in 1. It is also desired that contributions of different layers and filters to loss landscape are comparable, since otherwise some NN elements can overshadow all others. This would also allow to compare loss landscapes of NNs with different architectures. Another illustration of the necessity of normalization is related to the scale invariance of rectified NNs with ReLU activations [3, 13]. It has to do with the observation that if weights of one layer are scaled up (multiplied) by some amount and weights of the next layer are scaled down (divided) by the same amount, the output of the neural network will not change. However, architectures with and without scaled weights will produce different loss landscapes, making our visualization somewhat arbitrary.
Numerically, normalization is concerned with adjusting the values in direction vectors depending on weights of the studied neural network. There are four normalization methods which include weight [16], filter [3], layer [3], and model [7] normalization methods, all of which are realized in LLA. The first one is the most intuitive yet fails to account for scale invariance. The latter two normalize the direction of layer and model vectors, but do not consider individual filters. Filter normalization proposed in [3] seemingly satisfies all requirements of a good normalization method. However, in this paper we report that filter normalization, along with other method, does not prevent “value explosion” observed for some NNs. This is because while normalization prevents the uncontrollable growth of weights, it does not prevent layer outputs from changing uncontrollably. This ultimately means that the necessity to propose an optimal normalization method still exists.
2.3 Batch norm layers and “value explosion” in loss landscapes
Considering the discussion above, one might expect BN to be a natural solution since it normalizes the outputs of intermediate layers. However, one first has to consider that BN layers behave [17] differently in train and eval operating modes of neural networks [18]. BN statistics are calculated for every input batch and used to learn BN parameters in train mode, which are later applied during inference in eval mode. Therefore, layer output normalization is guaranteed only in train mode, whereas the results in eval mode depend on the similarity between training and evaluation data.
Furthermore, when we modify weights to plot loss landscapes, we potentially make learned BN weights irrelevant. This behavior was previously observed by other researches who proposed to not modify BN weights with direction vectors [3]. However, in our experiments this approach did not yield any positive results, which can partly be due to LLA plotting loss landscapes for a rather vast region surrounding the point corresponding to the original weights. The most robust solution is plotting loss landscapes in train mode allowing BN layers to work as intended preventing value explosions. This is a problem since eval mode is the main inference mode of neural networks used in all applications “in the Wild”.
That being said, one could argue that such effects of BN on loss landscapes is a mere artifact of the method. In real inference scenarios the changes in dataset features cannot really lead to NN value explosion due to preprocessing, input data normalization, etc. Hence, value explosion in loss landscapes has to do only with the way we alter the weights in (1), and using BN in train mode allows NN to adjust to these changes similar to how preprocessing works for unseen data. Therefore, using train mode gives results that are actually more meaningful for the analysis of inference modes. However, this argument does not resolve the ambiguity between train and eval mode plotting completely, illustrating the need for a new normalization method.
2.4 Loss landscapes of popular NNs
In this Section we examine loss landscapes of popular networks to investigate the applicability of the existing visualization methods. These are obtained for randomly initialized and ImageNet-1K pre-trained NNs. All models and their pre-trained weights are taken from PyTorch library [19] with the exception of VIT [20].
Table 1 shows that value explosion is observed for most CNNs in various regimes. Figure 2 shows that when value explosion occurs, no meaningful information can be obtained unless a regime which avoids it is chosen. In addition to train mode, changing axes from random to Hessian sometimes solves this problem. Table 1 also shows that there is no normalization mode that allows to avoid value explosion completely. Furthermore, filter normalization always leads to value explosion if the NN is prone to it. Using L1 instead of L2 norm in filter normalization can sometimes help to avoid this.
It should be noted that MobileNet and VIT whose landscapes are shown in Figure 3 are the only studied NNs that do not exhibit value explosion in any mode. Loss landscapes of AlexNet [21], SqueezeNet [22], and LeNet [23] behave almost identically and have asymmetric minimum when value explosion occurs, which can be viewed when loss is capped at some value, as shown in Figure 4.
Table 1: The behavior of loss values in loss landscapes of CNNs and VIT; digits refer to normalization methods: 1 – none, 2 – weight – 3 filter L1, 4 – filter L2.
model Explosion in train, Explosion in train, Explosion in eval, Explosion in eval,
random axes Hessian axes random axes Hessian axes
LeNet yes134 no234 yes134 no234
ResNet no no yes yes234
AlexNet yes134 yes24 yes134 yes24
SqueezeNet yes124 yes24 yes124 yes24
MobileNet no no no no
VIT no no no no
yesn…m means that only normalization methods n…m lead to value explosion,
non…m means that for normalization methods n…m values do increase significantly (but do not explode).
Figure 2: Loss landscapes for ResNet18 in different regimes: (left) value explosion in eval mode, (center) normal behavior with random axes in train mode, (right) normal behavior with hessian axes in train mode.
Figure 3: (Left) MobileNet and (right) VIT loss landscapes with filter L2 normalization which exhibit no “value explosion”.
Figure 4: Filter-normalized loss landscapes of models that exhibit value explosion, typical for SqueezeNet, AlexNet, and LeNet: (left) no loss cap and (right) loss caped at 100.
3 Hessian of neural networks and Hessian analysis criteria
In context of NNs, Hessian is a matrix of second derivatives of loss function with respect to weights that incorporates all information about the curvature of loss function at a point. Most importantly, Hessian eigenvalues can provide valuable information about local curvature of loss landscapes. Since forming Hessian explicitly is a very costly task, the applicability of Hessian analysis and other second order methods is somewhat limited. However, owing to the recent advances in Randomized Numeric Linear Algebra (RandNLA) [24, 25], properties of the Hessian which include its eigenvalues, trace, and eigenvalue spectral density can be evaluated stochastically even when Hessian matrix is not available [26, 27]. This facilitates its application to neural network analysis problems.
3.1 Hessian axes and spectral density
Conventionally, Hessian eigenvalues and trace have been the most common criteria for Hessian analysis [13, 27]. It was observed that eigenvalues and trace grow for poorly trained networks making their values a possible indicator of network’s performance. However, absolute values of these parameters depend on specifics of NN architectures and weight values, which, as has been discussed in Section 3.1, do not necessarily have to be unique for NNs to perform identically. This makes the use of eigenvalues and trace somewhat unreliable especially when different networks are to be compared.
However, there is a very clear geometric relation between Hessian eigenvalues and loss landscape curvature. That is, negative eigenvalues are indicative of the presence of areas with negative curvature, whereas positive ones correspond to the areas of positive curvature. That inspired some researchers to propose the ratio of negative and positive eigenvalues to be a measure of local curvature of loss landscapes [3] which unfortunately requires evaluating Hessian eigenvalues at every point of the landscape. Another approach is the evaluation of the full Hessian Eigenvalues Spectral Density (HESD), which provides information about all Hessian eigenvalues [27]. This is useful since the sole presence of large eigenvalues may give limited information in case that the “weight” of this eigenvalue (and the direction of associated eigenvector) is relatively small in the spectrum.
In this paper we mainly focus on HESD evaluation as the main tool for Hessian analysis. It is based on Stochastic Lanczos Algorithm and for the exact derivation the reader is referred to [26, 27, 28]. LLA also provides additional useful applications for Hessian eigenvectors, as they can be used as deterministic direction vectors for weight update in (1). Specifically, LLA uses two eigenvectors that correspond to two largest Hessian eigenvalues as Hessian directions. All results which use “Hessian axes” in this paper have been obtained using this method. This method is similar to the one proposed in [14] and allows to obtain useful results while avoiding some drawbacks of random direction method, as Section 2 showed.
3.2 Typical HESD types
When researching HESD application to NN analysis, one might find similar HESDs for different NN architectures reported by different research teams [13, 27, 29]. This observation inspired us to conduct an investigation into HESDs of different networks that concluded that their structure is almost universal. This Section aims to list typical HESDs and provide an explanation to this phenomenon.
The first observation is that HESD plots of untrained NNs with weights randomly chosen from a uniform distribution are symmetrical with respect to zero. Figure 5 illustrates this for several CNNs, and this behavior has also been observed for much larger class of NNs. Whereas the exact shape of HESD is defined by network architecture, the symmetry is a consequence of symmetrical structure of weight matrices which, when filled with uniformly distributed random numbers, yields no preferred direction in Hessian space resulting in a symmetric spectrum.
Figure 5: HESD plots of untrained neural networks with randomly initialized weights: (left) AlexNet, (center) SqueezeNet 1.1, and (right) MobileNetV2.
The second observation is that a fully trained neural network with very low training loss and high training accuracy has almost exclusively positive eigenvalues, as shown in Figure 6. This can be explained by the fact that during training optimizers aim to change model weights along the negative slope of the loss landscape. Hence, for a fully trained NN there is essentially no negative slope left, which is reflected in the near absence of negative eigenvalues in HESD.
Figure 6: HESD plots of neural networks trained to over 99% accuracy: (left) LeNet trained on MNIST and (right) ResNet20 trained on Cinic10.
Finally, in intermediate states between random initialization and being fully trained, i.e., during training, the negative eigenvalue section of HESD reduces while the positive section grows. This is illustrated by Figure 7 and other results obtained in this study. This behavior can also be explained by the gradual reduction in negative eigenvalues during training accompanied by gradual emergence of large positive eigenvalues. The latter were previously referred to as outliers which correspond to high positive curvature directions associated with classifier classes of studied NNs [13]. Near-zero eigenvalue sections also correspond to a large number of degenerate directions in Hessian space which indicates a mostly flat loss landscape.
Figure 7: HESD plots of neural networks trained to 70-80% accuracy: (left) SqueezeNet 1.1 pre-trained on ImageNet-1K, (center) MobileNetV2 pre-trained on ImageNet-1K, and (right) LeNet trained on MNIST.
Such HESD plots have been observed for a large class of convolutional and fully-connected NNs. Similar HESD plots were also observed for Generative Pretrained Transformers (GPTs) [29]. It was suggested that typical HESD structures can be attributed to block structure of NNs and the same loss function (cross-entropy) used for all classifiers, which also holds true for our results.
However, two exceptions to the above behavior have been observed. First, ResNets have asymmetric HESD plot upon random initialization, as shown in Figure 8. This effect occurs only in train mode of the neural network when BN parameters are calculated using input batch data statistics, and it is not present for evaluation mode when all parameters are fixed. This means that the asymmetry which arises from parameter calculation in BN layers is sufficient to create relatively large negative eigenvalue sections in HESD. However, this effect is either minor or not present in other NNs with BN layers.
Figure 8: HESD plots of randomly initialized ResNet18 in (left) train and (center) eval modes, and (right) its ImageNet-1K pre-trained version.
Lastly, whereas HESD of randomly initialized VIT behaves as expected, HESD of pre-trained VIT shown in Figure 9 is different from the previously discussed ones. It is characterized by a large negative eigenvalue section, even for VITs that have relatively high training accuracy and low loss. This is also peculiar since VIT loss landscape shown in Figure 3 does not show any large negative curvature. Whereas this observation is very interesting, it requires additional study and the detailed analysis of VIT behavior will be conducted later.
Figure 9: HESD plot of (left) randomly initialized and (right) ImageNet-1K pre-trained VIT.
3.3 Hessian analysis criteria
In this Section we propose several criteria that can be conveniently calculated using HESD to analyze it without necessarily having to visually inspect the plots. Following the discussion in Section 3.1, the first criterion is simply the ratio of most negative and most positive Hessian eigenvalues
$r_{\text{e}}=\frac{\max\left(\mathrm{abs}\left(\lambda_{\text{neg}}\right)\right)}{\max\left(\lambda_{\text{pos}}\right)},$ (2)
where $\lambda_{\text{neg}}$ and $\lambda_{\text{pos}}$ are negative and positive Hessian eigenvalues, respectively. However, this criterion does not take into account the complete eigenvalue spectrum. To address this issue, the following metric is proposed
$K_{\text{Hn}}=\frac{\sum\limits_{i}\left(\mathrm{abs}\left(\lambda_{\text{neg,i}}\right)\cdot w_{\text{neg,i}}\right)^{n}}{\sum\limits_{j}\left(\lambda_{\text{pos,j}}\cdot w_{\text{pos,j}}\right)^{n}},$ (3)
where $\omega_{\text{neg}}$ and $\omega_{\text{pos}}$ are weights corresponding to negative and positive eigenvalues used to form HESD (see Section 3C in [27] for details), respectively, and n is some positive real number.
To investigate whether these criteria reflect the changes in HESD and NN performance correctly, we calculate re, KH1 (n = 1 in (3)), and KH05 (n = 0.5 in (3)) for several neural networks and compare their random initialization states to the pre-trained ones, as shown in Table 2. ImageNet-1K pre-trained weights were used for all neural networks but LeNet, which was trained on MNIST as part of this study. Accuracy is calculated for the same batch of training data that was used to evaluate HESD of pre-trained models and averaged for train and eval experiments.
Table 2: HESD criteria for randomly initialized and pre-trained neural networks. Three values in cells correspond to KH1 , KH05 , and re, respectively.
model Random init, Pre-trained, Random init, Pre-trained, Average
train mode train mode eval mode eval mode accuracy
Lenet 0.99, 0.95, 0.88 0.37, 0.23, 5e-4 0.98, 0.98, 0.8 0.18, 0.17, 3e-4 99 %
ResNet18 1.15, 2.3, 451 1.03, 0.67, 0.32 0.97, 0.87, 0.73 0.97, 0.66, 0.03 86 %
AlexNet 1.0, 1.0, 1.0 0.94, 0.64, 0.04 0.99, 0.97, 1.0 0.97, 0.67, 0.06 84 %
SqueezeNet 1.18, 1.23, 0.96 0.7, 0.43, 0.02 1.08, 1.23, 1.01 0.76, 0.55, 0.02 75 %
MobileNet 1.07, 1.21, 5.7 1, 0.75, 0.1 1.04, 1.0, 1.0 0.83, 0.54, 0.01 84 %
VIT 0.99, 0.98, 0.9 1.03, 1.21, 1.85 0.99, 0.95, 0.91 0.99, 1.04, 1.87 87 %
Table 2 shows that whereas KH1 works for LeNet trained to very high accuracy, it fails to capture the changes between pre-trained and randomly initialized states of other CNNs. On the contrary, KH05 consistently decreases when pre-trained weights are used. That being said, there is no direct correlation between KH05 and accuracy values, since for networks with 85% average accuracy KH05 ranges from 0.64 to 0.75. Therefore, it is the change in KH05 value but not its exact value that can be used to access neural network’s performance. It is also important that it allows to obtain some estimate of NN performance while requiring only a few (even just one) batches of input data for its calculation.
Table 2 also shows that re decreases for pre-trained models, too. However, its variation range is extremely large making it less convenient to use as a criterion compared to KH. It also sometimes does not reflect the changes in accuracy correctly, as will be shown in Section 4. On the other hand, it can help to capture abnormalities like the behavior of randomly initialized ResNet18 in train mode, for which re sky-rockets to 451. It should also be noted that the strange behavior of VIT HESD discussed previously results in all criteria giving opposite results for VIT compared to CNNs.
3.4 Computational speed and stability of HESD criteria
Since HESD is calculated using stochastic methods, it can vary depending on calculation seed. Our tests have shown that Lanczos part of HESD evaluation algorithm is, predictably, the source of results’ variations. Therefore, the criteria proposed in the previous Section can also vary, which brings their reliability into question.
It was determined that weights in (3) are affected by stochastic nature of the algorithm more than eigenvalues in (2) and (3), making re a more computationally stable criterion compared to KH. Regarding the latter, several experiments were conducted to investigate the optimal conditions for KH application. NVIDIA A100 GPU was used for all experiments.
First, the dependence of the results on the number of HESD evaluation runs (n_hes) was studied. When running consequent calculations with different seeds, for instance, for average KH05 = 0.16 with variation is 0.13-0.2. n_hes = 10 allows to get relatively stable results which vary slightly with later calculations. Increasing n_hes to 20 allows to obtain more reliable results, though it is more computationally expensive.
Another aspect that is not connected with stochastic uncertainties but can affect the results is the choice of data for HESD calculation. While one batch of data is sufficient to calculate the criteria, the results may vary for different batches in the dataset. It was noticed that even for high n_hes the variation can be substantial, for instance, for average KH05 = 0.16 the variation is 0.1-0.24 for LeNet trained on MNIST. However, averaging for N random batches allows to get relatively stable results. For the same example with KH05 = 0.16, the variation drops to 0.05 when at least N = 4 batches are used. Similar results were observed for other studied cases, including ResNet(ImageNet) experiments. Using several batches also allows to use fewer n_hes.
However, the necessity to conduct multiple calculations to stabilize the results questions the efficiency of the proposed method in comparison to simple accuracy evaluation. For instance, for LeNet on MNIST the criterion calculation on four batches of 2048 images with n_hes = 10 takes ten times more time than accuracy calculation on full dataset. This shows that the proposed criteria are suboptimal for small dataset and compact networks. On the contrary, for large models and extremely large datasets like ResNets(ImageNet-1K) the advantages of the proposed criteria become apparent, since KH calculation takes roughly 22 sec for four 64 image batches, which is equivalent to accuracy calculation on 3500 out of several million ImageNet images [30].
4 Generalization experiments
In this Section we investigate how loss landscape and Hessian analyses can be used to assess the generalization capabilities of NNs. In order to achieve this, we use two pairs of datasets with compatible classes to conduct generalization experiments. This allows NNs trained on one dataset be used for inference on another one without retraining or fine-tuning. This setup is more similar to real generalization scenarios than the setups that use test subset of training dataset to calculate generalization accuracy [3].
4.1 Datasets
Two pairs of datasets were used to conduct generalization experiments: digits’ datasets MNIST [31] and “The Street View House Numbers” (SVHN) [32], and image datasets Cifar10 [33] and Cinic10 [34]. All datasets have 10 classes with 32x32 images, which are grayscale for MNIST and RGB for others. SVHN images were converted to grayscale to be compatible with MNIST.
It should be noted that Cinic10 consists of Cifar10 and some ImageNet images relabeled with Cifar10 class labels. In order to make generalization experiments valid, all Cifar10 images have been removed from Cinic10 with this version being referred to as Cinic10-i in Table 3. Validation split of Cinic10 has not been included into Cinic10-i. For simplicity we will refer to Cifar10 as cifar and to Cinic10-i as cinic in this paper. All datasets and their train/test splits are described in Table 3.
Table 3: Description of the datasets used in this study.
Dataset Image shape Dataset size Train split Test split Note
size size
MNIST 32x32x1 70000 60000 10000 -
SVHN 32x32x1 99289 73257 26032 Converted to grayscale
from original RGB.
Cifar10 32x32x3 60000 50000 10000 -
Cinic10 32x32x3 270000 90000 90000 Cifar10 and ImageNet
images
Cinic10-i 32x32x3 140000 70000 70000 ImageNet images only
In addition, one batch of 64 images was collected from the internet and manually labeled with ImageNet-1K labeles to test generalization capabilities of popular models pre-trained on ImageNet. To minimize the odds that some of the images were part of ImageNet training subset, only relatively recent images not used on popular websites were considered. Images were rescaled to 224x224x3 to match the input size requirements of the pre-trained models. This dataset is referred to as NI in Section 4.4.
In all following experiments models are trained for 1000 epochs with 0.001 learning rate using Adam optimizer on the train split of the training dataset which is used for training accuracy calculation. When reporting results, training dataset is indicated in brackets after NN name, e.g. MNIST in LeNet(MNIST). Generalization experiments consist of conducting model inference on the test subset of another dataset, and then calculating generalization accuracy on that subset. This dataset is indicated after dash symbol, e.g. SVHN in LeNet(MNIST)-SVHN. The experiments constituting of NN training and testing on the same dataset are not included to avoid confusion.
4.2 LeNet MNIST-SVHN generalization experiments
The first dataset pair was used to train LeNet. Figure 10 shows that generalization accuracy is much lower when generalizing from MNIST to SVHN. This behavior can be expected since a smaller dataset is used for training. LeNet(MNIST)-SVHN also shows abnormal growth in generalization accuracy after the training is essentially finished, which is not observed in other experiments. However, this increase amounts only to several percent not even reaching 20% generalization accuracy at 1000th epoch.
Figure 10: Training loss, training accuracy, and generalization accuracy (left) LeNet(MNIST)-SVHN and (right) LeNet(SVHN)-MNIST.
On the contrary, LeNet exhibits 55-58% generalization accuracy on MNIST when trained on SVHN. It should be noted that higher generalization accuracy is accompanied by lower training accuracy that does not reach 90%. In both cases loss continues to decrease even after maximum accuracy is reached, and for LeNet(SVHN) experiment this does not have any significant effect on generalization accuracy.
As has been previously shown in Table 1, LeNet loss landscape exhibits value explosion in many regimes which limits the applicability of loss landscape analysis to its generalization capability evaluation. Figure A.1 shows loss landscapes for LeNet(MNIST) with weight normalization using random axes that show drastic increase in loss value when tested on SVHN. However, loss values seem to generally be lower for MNIST, since for generalization landscape LeNet(SVHN)-MNIST loss values decrease, too. This makes loss landscape analysis in this case ambiguous because no direct correlation between landscape changes and generalization accuracy can be derived.
However, Figure 11 shows that KH05 increases significantly when changing datasets. Furthermore, this correlates with changes in generalization accuracy, since KH05 increases (and accuracy drops) for MNIST-SVHN case much more drastically, than for SVHN-MNIST case. Plots for other criteria are shown in Appendix B.
Figure 11: The results for KH05 for (left) LeNet(MNIST)-SVHN and (right) LeNet(SVHN)-MNIST.
4.3 ResNet20 cifar-cinic generalization experiments
Figure 12 shows that the final training accuracy of approximately 97% is reached around 200th epoch for both ResNet20(cinic) and ResNet20(cifar). However, generalization accuracy reached its peak value at approximately 70th epoch and then decreased to around 43% maintaining this value with roughly 1% oscillation for both experiments. The highest generalization accuracy of 55% is reached by ResNet(cinic)-cifar, while ResNet(cifar)-cinic genealization accuracy does not exceed 45%.
It is interesting that unlike MNIST-SVHN experiment where training on smaller dataset results in much worse generalization accuracy, ResNet20(cifar)-cinic reaches roughly the same training and generalization accuracies as ResNet20(cinic)-cifar. This might be explained by very similar feature distributions in both datasets [34].
Figure 12: Training loss, training accuracy, and generalization accuracy of (left) ResNet20(cinic)-cifar and (right) ResNet20(cifar)-cinic.
Figure A.2 shows that loss landscapes of ResNet20 are overall rather non-uniform even when plotted for training data. Therefore, even that the cifar-cinic plot is clearly more chaotic than the cifar training one, it is hard to give quantitative estimates to the changes in NN performance. On the contrary, the results for criterion KH05 shown in Figure 13 indicate that it increases in generalization experiments, thus corresponding to the decrease in accuracy. However, the re criterion is smaller for generalization than re for training, not reflecting the accuracy changes correctly, as figures in Appendix B show.
Figure 13: The results for criterion KH05 for (left) ResNet20(cinic)-cifar and (right) ResNet20(cifar)-cinic.
This means that the increase in KH05 when changing datasets for a trained model can indicate poor generalization and the reduction in accuracy. However, as mentioned before, there is no strict correlation between KH05 and accuracy values. Indeed, the average KH05 value for ResNet20(cinic)-cifar is 0.25, and its value for ResNet20(cifar)-cinic is 0.45, whereas both correspond to approximately 43% generalization accuracy.
4.4 Generalization capability of popular NNs
Table 4 summarizes the results of generalization experiments conducted with several popular CNNs and VIT. One random batch of 64 ImageNet-1K images is used to calculate ImageNet HESD criteria and accuracy. In all cases ImageNet-1K pre-trained weights were used for model inference.
Table 4: HESD criteria and inference accuracy of popular CNNs and VIT.
model dataset Accuracy, % KH1 KH05 re
ResNet50 ImageNet 85.9 0.86 0.48 0.064
ResNet50 NI 71.9 0.95 0.65 0.065
AlexNet ImageNet 84.3 0.94 0.65 0.11
AlexNet NI 23.4 0.96 0.8 0.21
SqueezeNet ImageNet 75 0.72 0.5 0.06
SqueezeNet NI 23.4 0.88 0.64 0.086
MobileNet ImageNet 84.3 0.71 0.40 0.015
MobileNet NI 59.3 0.76 0.54 0.04
VIT ImageNet 87.5 0.99 0.96 6.3
VIT NI 89.1 0.95 0.82 0.77
Table 4 shows that KH05 behaves as has been previously discussed in this study, increasing when generalization accuracy drops for CNNs. As mentioned earlier, KH1 does not always have this property, as AlexNet results show. In many cases re also increases when accuracy decreases, though these changes vary significantly for different experiments, and re behavior does not always reflect that of the accuracy. This makes KH05 the most suitable Hessian criterion for NN generalization capability assessment.
However, VIT is once again an exception. It should also be noted that NI accuracy being higher than ImageNet accuracy for VIT can be explained by the fact that the class set in NI is very unambiguous, resulting in a simpler tasks compared to the analysis of random ImageNet batches.
5 Conclusions
This paper discusses the applicability of loss landscape and Hessian analyses to the study of generalization capabilities of NNs while focusing on different aspects of the methodology aiming to propose new analysis criteria. The results were obtained using a Loss Landscape Analysis library developed as part of this study. Different approaches to loss landscape evaluation were discussed showing that conventional methods might fail in specific circumstances which raises the demand for an improved normalization method. This conclusion was made after a comprehensive analysis of loss landscapes of conventional NNs in various regimes was conducted. It was shown that using Hessian axes can sometimes mitigate this issue, and the approach for choosing Hessian axes was proposed. Typical Hessian spectra were also discussed and several criteria for their analysis were proposed, while highlighting the differences between CNNs and VIT. Generalization experiments conducted using LeNet and ResNet20 on MNIST-SVHN and cifar-cinic dataset pairs showed that the proposed criteria can be used as an estimate of generalization accuracy, which is computationally efficient for large datasets.
Acknowledgement
The author would like to thank his colleagues Dr Anton Raskovalov, Dr Igor Netay, and Ilya Androsov for fruitful discussions, and Vasily Dolmatov for discussions and project supervision.
Appendix A Appendix A
Figure A.1: Random axes weight-normalized loss landscapes of LeNet (top-left) trained and tested on MNIST, (top-right) trained on MNIST and tested on SVHN, (bottom-left) trained and tested on SVHN, and (bottom-right) trained on SVHN and tested on MNIST. All results correspond to 249th epoch of respective experiments.
Figure A.2: Random axes filter L2-normalized loss landscapes of ResNet20 (top-left) trained and tested on cinic, (top-right) trained on cinic and tested on cifar, (bottom-left) trained and tested on cifar, and (bottom-right) trained on cifar and tested on cinic. All results correspond to 249th epoch of respective experiments.
Appendix B Appendix B
Figure B.1: Hessian criteria for LeNet(MNIST)-SVHN: (left) KH1, (center) re from max to min values, and (right) re in near-zero area.
Figure B.2: Hessian criteria for LeNet(SVHN)-MNIST: (left) KH1, (center) re from max to min values, and (right) re in near-zero area.
Figure B.3: Hessian criteria for ResNet20(cinic)-cifar: (left) KH1, (center) re from max to min values, and (right) re in near-zero area.
Figure B.4: Hessian criteria for ResNet20(cifar)-cinic: (left) KH1, (center) re from max to min values, and (right) re in near-zero area.